# Superconducting pairing correlations on a trapped-ion quantum computer

### v1

@@ V Conclusion
First and foremost are the limits placed on adiabatic algorithms caused by entropy accumulation due to hardware errors. One option that appears sufficient (though maybe not necessary) to reach low-temperature states is to adiabatically evolve large systems for hundreds of hopping times [10]. This is likely achievable with a factor 1000 more Trotter steps than we explored here [39, 40, 41, 42], which translate to per-two-qubit-gate errors on the order of $10^{-6}$. While this requirement is mild compared to other quantum computing applications, understanding the most effective combination of (fermionic) error corrections codes [43, 44] and algorithmic techniques like bath-assisted cooling [45, 46] is essential to reach this goal. Reductions in gate error can also be used to simulate more complex Hamiltonians or longer evolution times which may be used to gain a better microscopic understanding of light-induced superconductivity in experimental setups [47, 48, 9].
On the technical side, we have introduced techniques to coherently inject states with locally fixed fermionic parity into fermionic encodings. These circuits can be used to access and explore superconducting phases and they show that ground state preparation for spin models can be useful beyond the study of magnetism. The techniques we have demonstrated can be adapted to other geometries, next-nearest neighbour hopping with a negative sign [6], and other Hamiltonian terms, imparting, for example, the ability to introduce dopants coherently.
@@
One difficulty in measuring pairing correlations is that their local fermionic parity is odd when restricted to a single spin species. Previous Hubbard model simulations on quantum computers have used encodings in which the up- and down-fermions are encoded separately [49, 50, 51, 25]. On the one hand, this makes it relatively easy to find efficient circuits for the time evolution under the hopping terms of the Hamiltonian $t\sum_{\langle ij\rangle,\sigma}c^{\dagger}_{i\sigma}c_{j\sigma}$. On the other hand, in such encodings, measuring distant pairing correlations requires extended Jordan-Wigner strings, which makes it difficult to simultaneously measure many correlators, and leads to high-weight operators, which are highly susceptible to noise [39].
@@
To overcome these difficulties, in Figure S2, we show an “Octagon” fermion-to-qubit encoding. The encoding rules are closely related to the Compact Encoding [22, 23] and are as follows: Each site $i$ of the square-lattice Hubbard model is encoded by two qubits, one representing the “up” (blue, $\sigma=\uparrow$) and one representing the “down” (red, $\sigma=\downarrow$) fermionic modes. A diagonal edge is placed between each such pair. The square lattice is partitioned into even and odd plaquettes, corresponding to a checkerboard pattern in Fig S2a, with the plaquette at the south-west corner being an odd plaquette. Then, the odd plaquettes are further partitioned into $P$-plaquettes and $Q$-plaquettes, as illustrated in Fig S2a, so that plaquettes of one type contain disjoint sets of qubits. Edges are now placed between all nearest-neighbour “up”-qubits in the $P$-plaquettes and between all nearest-neighbour “down”-qubits in the $Q$-plaquettes. Depending on their geometry, these edges are horizontal or vertical edges. Within each odd plaquette, an ancilla qubits is placed. Thus each horizontal and vertical edge $(i\sigma,j\sigma)$ has a unique ancilla $a(i,j)$ associated to it. Edges are assigned an orientation, as shown via the arrows in Figure S2. Finally, identification between fermionic modes and qubit operators is made by associating to each edge $(i\sigma,j\sigma^{\prime})$ the Majorana bilinear $A_{i\sigma,j\sigma^{\prime}}=-i\gamma_{i\sigma}\gamma_{j\sigma^{\prime}}$, where $\gamma_{i\sigma}=c^{\dagger}_{i\sigma}+c_{i\sigma}$ and the qubit operator $A_{i\sigma,j\sigma^{\prime}}=X_{i\sigma}Y_{j\sigma^{\prime}}P$, where $P$ is $Y_{a(i,j)}$ for horizontal edges, $X_{a(i,j)}$ for down-pointing vertical edges, $X_{a(i,j)}$ for up-pointing vertical edges, and $\mathds{1}$ for diagonal edges. Note that the order of the indices of $A_{i\sigma,j\sigma^{\prime}}$ follows the orientation of arrows on the lattice (in this case $i\sigma\rightarrow j\sigma^{\prime})$, and reversed operators are defined as $A_{j\sigma^{\prime},i\sigma}=-A_{i\sigma,j\sigma^{\prime}}$. To complete the algebra of fermionic bilinears, we associate to each mode $i\sigma$ an on-site term $B_{i\sigma}=-i\gamma_{i\sigma}\overline{\gamma_{i\sigma}}\rightarrow Z_{i\sigma}$, where $\overline{\gamma_{i\sigma}}=i(c^{\dagger}_{i\sigma}-c_{i\sigma})$. Edge operators between non-adjacent modes $(i\sigma,k\sigma^{\prime})$ can be defined if there exists a path of edges $\{(j_{1},j_{2}),\dots(j_{l-1},j_{l})\}$, via $A_{i,k}=i^{l-1}\prod_{m=1}^{l-1}A_{j_{m},j_{m+1}}$, where $j_{1}=i\sigma$ and $j_{l}=k\sigma^{\prime}$. For paths $j_{1},\dots j_{l}$ that form closed loops $j_{l}=j_{1}$, one associates an operator $S=iA_{j_{1},j_{l}}$. The eigenvalues of $S=\pm 1$ indicate whether or not there is a $\pi$-magnetic flux through the enclosed area. If the flux-free sector is desired, one needs to ensure $S=1$ for all closed loops (this is the case for all topologically trivial loops of the experiments in this work, except for the closed loops that wrap around the torus, which are set to $\langle S\rangle=0$ to obtain mixed boundary conditions, see Section S13). Note that, in the qubit representation, the constraints on minimal loops that are topologically trivial are given by equations of the form $Z^{\otimes 8}XYXY=+1$, whereas if periodic boundary conditions were desired then one would need to set the qubit operators for minimal topologically non-trivial loops on the $6\times 6$ lattice (Figure S2) $Z^{\otimes 12}Y^{\otimes 6}=-1$ and $Z^{\otimes 12}X^{\otimes 6}=-1$, i.e., there is a negative sign for the non-trivial operators. The difference in sign arises due to the fact that the topologically trivial elementary loops touch an odd number of up-pointing vertical edges (namely, one), which are associated with a X-X operator on the ancilla.
@@
Figure S3: Octagon Fermion-to-Qubit Encoding for the operators and observables in the bilayer Hubbard Model. (a) Hamiltonian terms in the bilayer Hubbard model that is used to model bilayer nickelates [35, 36, 37, 38]. The terms within each layer are the same as for in the single-layer case, namely a nearest-neighbour hopping $t\sum_{\langle ij\rangle\sigma}c^{\dagger}_{i\sigma}c_{j\sigma}$ and an on-site interaction $U\sum_{i}n_{i\uparrow}n_{i\downarrow}$. The layers are coupled via an exchange coupling $J\sum_{i}\mathbf{S}_{iA}\cdot\mathbf{S}_{iB}-n_{iA}n_{iB}/4$. (b) The system is mapped to two separate Octagon Encodings, one containing all “up”-modes of layer $B$ and all “down”-modes of layer AA (top) and one containing all “down”-modes of layer $B$ and all “up”-modes of layer AA (bottom). This is possible because all three Hamiltonian terms, as well as the rung-singlet-pairing observable $\Delta^{\dagger}_{AiBi}\Delta_{AjBj}$ have even fermionic parity in each subsystem.
@@
In Fig S6, we investigate the effect of noise on stabilisers and on local observables through numerical simulations. We simulate the circuit used for preparing the $(D_{\mathrm{Heisenberg}},D_{\mathrm{Hubbard}})=(1,1)$ state with a Pauli string decomposition, see Section S12.2. We study the effect of depolarizing noise on the circuit, at a fixed Pauli string truncation threshold $\epsilon=0.1/2^{15}$. We observe that both the $\eta$-pairing correlation and the average stabiliser expectation value undergo an exponential decay with the noise rate. This very smooth behaviour of expectation value with depolarizing noise is to be put in contrast with truncation error in classical numerical simulations techniques (see for example section S12). We see however that the stabilisers (which are weight $12$ Pauli strings) are damped significantly faster than the $\eta$-correlations (which are weight $4$). Such a strong effect of observable weight on noise sensitivity was observed and explained before [39]. We also compare these classical simulations with the values from the quantum computer for both $\eta$-pairing and the stabilisers. The hardware value is compatible with the classical depolarising noise simulations within two standard errors on the mean.
@@
with an overview of the strategy shown in Fig. S8. We first remind that the edges appearing in this sum are not the edges present in the representation of the fermionic encoding in Fig S4a. The edges appearing in (S22) denote the edges on the original $6\times 6$ square lattice, and link only fermions with identical spins. In Fig S4a, they correspond for up spin (blue qubits) to edges such as 010-1 and 060-6, but also to edges such as 6126-12 and 787-8, even though they are not linked by an edge in this fermionic encoding. For down spin (red qubits), they include for example 434443-44 and 434943-49, but also 444544-45 and 433743-37. They do not include edges between different spins that appear in the fermionic encoding, such as 0360-36.
@@
Figure S7: Circuit to implement the plaquette hopping operator $U_{\rm hop}^{\rm plaq}$. Top: Direct implementation of the mapped operators via Pauli gadgets. Bottom: Reduction of two-qubit gate count and depth using circuit identities. Final gates (black) merge fermionic SWAP gates with interaction gates. One Hubbard-Trotter step requires two applications of the circuit. The total two-qubit gate count taking into account both hopping (coloured) as well as fSWAP/interaction gates (black) is thus $2\times(15N/2+N)=17N$ which, for the $N=6\times 6$ lattice evaluates to 612 two-qubit gates. The two-qubit gate depth is 18, independent of system size.
@@
The terms appearing in (S22) do not all commute with each other, so that further Trotterization is required to implement the exponential of $H_{\rm hop}$. In order to explain our choice of Trotterization, let us denote a $\sigma$-spin plaquette” (with $\sigma=\uparrow,\downarrow$) a group of $4$ qubits with same spin located in a square pattern around an ancilla, either directly around it or not directly around it. Namely, there are exactly two plaquettes per ancilla, one with spin up and one with spin down qubits. For example in Fig S4 (a), the plaquette with up spin around ancilla $72$ is comprised of qubits $0,1,6,7$, and the plaquette with down spin around ancilla $72$ is comprised of qubits $36,37,42,43$. In section S2, we further distinguished two types of plaquettes: we call $P$-plaquettes those associated to ancillas on even rows (equivalently, ancillas with up-spin qubits directly around them), and $Q$-plaquettes those associated to ancillas on odd rows (equivalently, ancillas with down-spin qubits directly around them). For example in Fig S4 (a), the plaquette $0,1,6,7$ is a up-spin $P$-plaquette, the plaquette $36,37,42,43$ is an down-spin $P$-plaquette, the plaquette $43,44,49,50$ is an down-spin $Q$-plaquette and the plaquette $7,8,13,14$ is a up-spin $Q$-plaquette.
@@
Let us consider one arbitrary fixed-spin plaquette, such as for example qubits 0, 11, 66, 77 in Fig S4 (a). The hopping terms on a fixed-spin plaquette can be implemented exactly without Trotter error through the following decomposition
@@
Note that this cancellation only occurs when simulating either both vertical or both horizontal terms first — alternating horizontal and vertical terms would lead to cross-plaquette currents in the commutator expansion. Further, we note that the hopping terms of plaquettes of both the same type ($P$ or $Q$) and same spin always commute, because they apply on disjoint sets of qubits. Similarly, plaquettes that have both a different type and a different spin always commute. We Trotterize the entire hopping Hamiltonian by applying first the hoppings on all the up-spin P-plaquettes, then on all the down-spin $Q$-plaquettes, then on all the up-spin $Q$-plaquettes, and then on all the down-spin $P$-plaquettes. Namely
$U_{\rm hop}=\underbrace{\prod_{\begin{subarray}{c}{\rm down-spin}\\ {\rm P-plaquette}\end{subarray}}U_{\rm hop}^{\rm plaq}\prod_{\begin{subarray}{c}{\rm up-spin}\\ {\rm Q-plaquette}\end{subarray}}U_{\rm hop}^{\rm plaq}}_{=:\exp(-itA)}\underbrace{\prod_{\begin{subarray}{c}{\rm down-spin}\\ {\rm Q-plaquette}\end{subarray}}U_{\rm hop}^{\rm plaq}\prod_{\begin{subarray}{c}{\rm up-spin}\\ {\rm P-plaquette}\end{subarray}}U_{\rm hop}^{\rm plaq}}_{=:\exp(-itB)}\,$, (S24)
@@
where $i^{\prime}$ and $j^{\prime}$ denote the qubits on the same physical sites as $i$ and $j$, but with opposite spins. For example, if $i=36$, then $i^{\prime}=0$; if $i=7$ then $i^{\prime}=43$. We note that in this expression, the hopping term in the middle is applied on the qubits $i,j$ (that are not linked by an edge in the fermionic encoding), not on the qubits $i^{\prime},j^{\prime}$ that are linked by an edge in the fermionic encoding. Gathering everything, we obtain the following expression
@@
The main physical effect of this Trotterisation scheme is thus to break the full translation invariance to a symmetry under translation by two sites. Furthermore, the second-order Floquet Hamiltonian of the hopping Trotterisation (as well as the term that is dominant for strong interactions $U\gg t$) is shown in Fig. S10. It is noteworthy that part of the double-commutator recovers the original hopping Hamiltonian $H_{\mathrm{hop}}=A+B$. That commutator thus does not lead to an undesired term in the Floquet Hamiltonian, but merely renormalises the hopping $\sim(1+\tau^{2})$ which could in principle be accounted for when targeting a specific evolution time (although we have not made use of this fact for the circuits in this work).
@@
Note that this is a geometrically non-local circuit which allows us to prepare the toric code on the $3\times 3$ periodic ancilla lattice using a two-qubit depth of 4. This is the reason that we are not using a scheme based on measurement and feed-forward: While such schemes offer circuit depths that are constant in system size, for this particular system size, the minimal depth of 4 can be achieved unitarily with a smaller number of overall gates  [52, 53].
@@
this time with $\delta=-2/3$ exactly. The discussion of the bilayer perturbation theory is similar to the one in [3], although in that reference an exchange coupling between two layers of tJt-J-models was considered, whereas here we allow for the perturbation $V$ to create virtual states with doubly-occupied sites and only turn on the on-site interaction during the adiabatic evolution later on.
@@
The objective function to maximize is the overlap between the target state and the state generated by the circuit. The optimization over the circuit parameters is carried out through gradient descent over the infidelity. Specifically, we utilize the quasi-second order gradient descent method, L-BFGS algorithm, from SciPy [57]. With the 10-layer circuit, we obtain the fidelity $\mathcal{F}=0.99478$ with the best circuit realization from 30 random initializations. Notice each general $U(1)$ symmetry-preserving gate requires 3 elementary entangling 2-qubit gates, e.g., ZZPhase on our hardware. Due to a cancellation at the first layer, the total depth in terms of elemantary entangling 2-qubit gate of the XXZ state preparation circuit is 29. The parameters for all $10\times 4=40$ general $U(1)$-preserving gates used in the quantum experiments are given in Listing 7.
@@
To carry out the optimization of finding parameters of the fermionic circuits, here we utilize the Jordan-Wigner transformation for simplicity and efficiency. The Jordan-Wigner ordering we choose follows the qubit indices ordering. We first obtained both the $|s\rangle$ and $|d\rangle$ states with the Jordan-Wigner transformation. Then, we transform the variational fermionic circuit correspondingly according to the Jordan-Wigner ordering into a variational circuit over qubits by inserting fSWAPs. That is for any fermionic gate acting over not adjacent fermionic modes in the Jordan-Wigner ordering, we insert the fSWAPs. For example, the $U^{f}_{(1,3)}=\textrm{fSWAP}_{(1,2)}U_{(2,3)}\textrm{fSWAP}_{(1,2)}$. For a fermionic gate acting over adjacent fermionic modes, we apply no additional transformation. Now the problem reduces back to the standard variational quantum circuit optimization. We optimize the circuit with the quasi-second order gradient descent method, L-BFGS algorithm, from SciPy [57] over 30 different random initialization and select the best result. The best average fidelity is (2C)/2=0.99446(2-C)/2=0.99446.
@@
We use exact state vector simulation to obtain the exact energy gradient of all parameters. The optimization is carried out with the quasi-second order gradient descent method, L-BFGS algorithm, from SciPy [57] over 30 random initializations.

### v2

@@ V Conclusion
First and foremost are the limits placed on adiabatic algorithms caused by entropy accumulation due to hardware errors. One option that appears sufficient (though maybe not necessary) to reach low-temperature states is to adiabatically evolve large systems for hundreds of hopping times [xu_neutral-atom_2025]. This is likely achievable with a factor 1000 more Trotter steps than we explored here [granet_dilution_2025, heyl_quantum_2019, mizuta_trotterization_2025, kovalsky_self-healing_2023], which translate to per-two-qubit-gate errors on the order of $10^{-6}$. While this requirement is mild compared to other quantum computing applications, understanding the most effective combination of (fermionic) error corrections codes [algaba_fermion--qubit_2025, chen_error-correcting_2024] and algorithmic techniques like bath-assisted cooling [matthies_programmable_2024, kishony_efficiently_2025] is essential to reach this goal. Reductions in gate error can also be used to simulate more complex Hamiltonians or longer evolution times which may be used to gain a better microscopic understanding of light-induced superconductivity in experimental setups [fausti_light-induced_2011, baykusheva_ultrafast_2022, fava_magnetic_2024].
On the technical side, we have introduced techniques to coherently inject states with locally fixed fermionic parity into fermionic encodings. These circuits can be used to access and explore superconducting phases and they show that ground state preparation for spin models can be useful beyond the study of magnetism. The techniques we have demonstrated can be adapted to other geometries, next-nearest neighbour hopping with a negative sign [xu_coexistence_2024], and other Hamiltonian terms, imparting, for example, the ability to introduce dopants coherently.
@@
One difficulty in measuring pairing correlations is that their local fermionic parity is odd when restricted to a single spin species. Previous Hubbard model simulations on quantum computers have used encodings in which the up- and down-fermions are encoded separately [arute_observation_2020, stanisic_observing_2022, hemery_measuring_2024, nigmatullin_experimental_2025]. On the one hand, this makes it relatively easy to find efficient circuits for the time evolution under the hopping terms of the Hamiltonian $t\sum_{\langle ij\rangle,\sigma}c^{\dagger}_{i\sigma}c_{j\sigma}$. On the other hand, in such encodings, measuring distant pairing correlations requires extended Jordan-Wigner strings, which makes it difficult to simultaneously measure many correlators, and leads to high-weight operators, which are highly susceptible to noise [granet_dilution_2025].
@@
To overcome these difficulties, in Figure S2, we show an “Octagon” fermion-to-qubit encoding. The encoding rules are closely related to the Compact Encoding [derby_compact_2021, derby2021compactfermionqubitmapping] and are as follows: Each site $i$ of the square-lattice Hubbard model is encoded by two qubits, one representing the “up” (blue, $\sigma=\uparrow$) and one representing the “down” (red, $\sigma=\downarrow$) fermionic modes. A diagonal edge is placed between each such pair. The square lattice is partitioned into even and odd plaquettes, corresponding to a checkerboard pattern in Fig S2a, with the plaquette at the south-west corner being an odd plaquette. Then, the odd plaquettes are further partitioned into $P$-plaquettes and $Q$-plaquettes, as illustrated in Fig S2a, so that plaquettes of one type contain disjoint sets of qubits. Edges are now placed between all nearest-neighbour “up”-qubits in the $P$-plaquettes and between all nearest-neighbour “down”-qubits in the $Q$-plaquettes. Depending on their geometry, these edges are horizontal or vertical edges. Within each odd plaquette, an ancilla qubits is placed. Thus each horizontal and vertical edge $(i\sigma,j\sigma)$ has a unique ancilla $a(i,j)$ associated to it. Edges are assigned an orientation, as shown via the arrows in Figure S2. Finally, identification between fermionic modes and qubit operators is made by associating to each edge $(i\sigma,j\sigma^{\prime})$ the Majorana bilinear $A_{i\sigma,j\sigma^{\prime}}=-i\gamma_{i\sigma}\gamma_{j\sigma^{\prime}}$, where $\gamma_{i\sigma}=c^{\dagger}_{i\sigma}+c_{i\sigma}$ and the qubit operator $A_{i\sigma,j\sigma^{\prime}}=X_{i\sigma}Y_{j\sigma^{\prime}}P$, where $P$ is $Y_{a(i,j)}$ for horizontal edges, $X_{a(i,j)}$ for down-pointing vertical edges, $X_{a(i,j)}$ for up-pointing vertical edges, and $\mathds{1}$ for diagonal edges. Note that the order of the indices of $A_{i\sigma,j\sigma^{\prime}}$ follows the orientation of arrows on the lattice (in this case $i\sigma\rightarrow j\sigma^{\prime})$, and reversed operators are defined as $A_{j\sigma^{\prime},i\sigma}=-A_{i\sigma,j\sigma^{\prime}}$. To complete the algebra of fermionic bilinears, we associate to each mode $i\sigma$ an on-site term $B_{i\sigma}=-i\gamma_{i\sigma}\overline{\gamma_{i\sigma}}\rightarrow Z_{i\sigma}$, where $\overline{\gamma_{i\sigma}}=i(c^{\dagger}_{i\sigma}-c_{i\sigma})$. Edge operators between non-adjacent modes $(i\sigma,k\sigma^{\prime})$ can be defined if there exists a path of edges $\{(j_{1},j_{2}),\dots(j_{l-1},j_{l})\}$, via $A_{i,k}=i^{l-1}\prod_{m=1}^{l-1}A_{j_{m},j_{m+1}}$, where $j_{1}=i\sigma$ and $j_{l}=k\sigma^{\prime}$. For paths $j_{1},\dots j_{l}$ that form closed loops $j_{l}=j_{1}$, one associates an operator $S=iA_{j_{1},j_{l}}$. The eigenvalues of $S=\pm 1$ indicate whether or not there is a $\pi$-magnetic flux through the enclosed area. If the flux-free sector is desired, one needs to ensure $S=1$ for all closed loops (this is the case for all topologically trivial loops of the experiments in this work, except for the closed loops that wrap around the torus, which are set to $\langle S\rangle=0$ to obtain mixed boundary conditions, see Section S13). Note that, in the qubit representation, the constraints on minimal loops that are topologically trivial are given by equations of the form $Z^{\otimes 8}XYXY=+1$, whereas if periodic boundary conditions were desired then one would need to set the qubit operators for minimal topologically non-trivial loops on the $6\times 6$ lattice (Figure S2) $Z^{\otimes 12}Y^{\otimes 6}=-1$ and $Z^{\otimes 12}X^{\otimes 6}=-1$, i.e., there is a negative sign for the non-trivial operators. The difference in sign arises due to the fact that the topologically trivial elementary loops touch an odd number of up-pointing vertical edges (namely, one), which are associated with a X-X operator on the ancilla.
@@
Figure S3: Octagon Fermion-to-Qubit Encoding for the operators and observables in the bilayer Hubbard Model. (a) Hamiltonian terms in the bilayer Hubbard model that is used to model bilayer nickelates [lu_interlayer-coupling-driven_2024, qu_bilayer_2024, schlomer_superconductivity_2024, oh_type-ii_2023]. The terms within each layer are the same as for in the single-layer case, namely a nearest-neighbour hopping $t\sum_{\langle ij\rangle\sigma}c^{\dagger}_{i\sigma}c_{j\sigma}$ and an on-site interaction $U\sum_{i}n_{i\uparrow}n_{i\downarrow}$. The layers are coupled via an exchange coupling $J\sum_{i}\mathbf{S}_{iA}\cdot\mathbf{S}_{iB}-n_{iA}n_{iB}/4$. (b) The system is mapped to two separate Octagon Encodings, one containing all “up”-modes of layer $B$ and all “down”-modes of layer AA (top) and one containing all “down”-modes of layer $B$ and all “up”-modes of layer AA (bottom). This is possible because all three Hamiltonian terms, as well as the rung-singlet-pairing observable $\Delta^{\dagger}_{AiBi}\Delta_{AjBj}$ have even fermionic parity in each subsystem.
@@
In Fig S6, we investigate the effect of noise on stabilisers and on local observables through numerical simulations. We simulate the circuit used for preparing the $(D_{\mathrm{Heisenberg}},D_{\mathrm{Hubbard}})=(1,1)$ state with a Pauli string decomposition, see Section S12.2. We study the effect of depolarizing noise on the circuit, at a fixed Pauli string truncation threshold $\epsilon=0.1/2^{15}$. We observe that both the $\eta$-pairing correlation and the average stabiliser expectation value undergo an exponential decay with the noise rate. This very smooth behaviour of expectation value with depolarizing noise is to be put in contrast with truncation error in classical numerical simulations techniques (see for example section S12). We see however that the stabilisers (which are weight $12$ Pauli strings) are damped significantly faster than the $\eta$-correlations (which are weight $4$). Such a strong effect of observable weight on noise sensitivity was observed and explained before [granet_dilution_2025]. We also compare these classical simulations with the values from the quantum computer for both $\eta$-pairing and the stabilisers. The hardware value is compatible with the classical depolarising noise simulations within two standard errors on the mean.
@@
The idea is shown in Fig. S7 and Fig. S8: The basic problem when Trotterising a sum of hopping terms on overlapping bonds is that an order must be chosen, i.e. one bond must be chosen before the other (alternatively particle number conservation can be broken [evered_probing_2025] which is unsuitable in this work since we rely on particle number conservation to measure pairing correlations). This breaking of time reversal-symmetry leads to error terms in the effective Hamiltonian that are currents from orbitals on the first bond to orbitals on the second bond e.g.,
[…]
$=c^{\dagger}_{i\uparrow}c_{k\uparrow}-c^{\dagger}_{k\uparrow}c_{i\uparrow}.$
However, by simulating both vertical before the horizontal terms within one plaquette,
[…]
$=0,$
we can cancel all commutators and implement hopping around $2\times 2$ plaquettes exactly. Let us now describe the exact sequence of gates and implementation in our chosen fermionic encoding in more detail.
@@
We first remind that the $\langle ij\rangle$ appearing in (S22) are not the edges present in the representation of the fermionic encoding in Fig S4a, but instead denote the edges on the original $6\times 6$ square lattice, and link only fermions with identical spins. In Fig S4a, they correspond for up spin (blue qubits) to edges such as 010-1 and 060-6, but also to edges such as 6126-12 and 787-8, even though they are not linked by an edge in this fermionic encoding. For down spin (red qubits), they include for example 434443-44 and 434943-49, but also 444544-45 and 433743-37. They do not include edges between different spins that appear in the fermionic encoding, such as 0360-36.
@@
Figure S8: Circuit to implement the hopping operator for one plaquette $U_{\rm hop}^{\rm plaq}=\exp(-it\sum_{\langle ij\rangle\in\mathrm{plaquette}}c^{\dagger}_{i\sigma}c_{j\sigma}+\mathrm{h.c.})$. Top: Direct implementation of the mapped operators via Pauli gadgets. Bottom: Reduction of two-qubit gate count and depth using circuit identities. Final gates (black) merge fermionic SWAP gates with interaction gates. One Hubbard-Trotter step requires two applications of the circuit. The total two-qubit gate count taking into account both hopping (coloured) as well as fSWAP/interaction gates (black) is thus $2\times(15N/2+N)=17N$ which, for the $N=6\times 6$ lattice evaluates to 612 two-qubit gates. The two-qubit gate depth is 18, independent of system size.
@@
The terms appearing in (S22) do not all commute with each other, so that further Trotterization is required to implement the exponential of $H_{\rm hop}$. In order to explain our choice of Trotterization, let us denote a $\sigma$-spin plaquette” (with $\sigma=\uparrow,\downarrow$) a group of $4$ qubits with same spin located in a square pattern around an ancilla, either directly around it or not directly around it. Namely, there are exactly two plaquettes per ancilla, one with spin up and one with spin down qubits. For example, in Fig. S4 (a), the plaquette with up spin around ancilla $72$ is comprised of qubits $0,1,6,7$, and the plaquette with down spin around ancilla $72$ is comprised of qubits $36,37,42,43$. In section S2, we further distinguished two types of plaquettes: we call $P$-plaquettes those associated to ancillas on even rows (equivalently, ancillas with up-spin qubits directly around them), and $Q$-plaquettes those associated to ancillas on odd rows (equivalently, ancillas with down-spin qubits directly around them). For example in Fig S4 (a), the plaquette $0,1,6,7$ is a up-spin $P$-plaquette, the plaquette $36,37,42,43$ is an down-spin $P$-plaquette, the plaquette $43,44,49,50$ is an down-spin $Q$-plaquette and the plaquette $7,8,13,14$ is a up-spin $Q$-plaquette. Note that each bond of the (periodic) square lattice is associated to one unique $P$-type or $Q$-type plaquette.
@@
Let us consider one arbitrary fixed-spin plaquette, such as for example qubits 0, 11, 66, 77 in Fig S4 (a). As discussed above, the hopping terms on a fixed-spin plaquette can be implemented exactly without Trotter error through the following decomposition
@@
Further, we note that the hopping terms of plaquettes of both the same type ($P$ or $Q$) and same spin always commute, because they apply on disjoint sets of qubits. Similarly, plaquettes that have both a different type and a different spin always commute. We Trotterize the entire hopping Hamiltonian by applying first the hoppings on all the up-spin $P$-plaquettes, then on all the down-spin $Q$-plaquettes, then on all the up-spin $Q$-plaquettes, and then on all the down-spin $P$-plaquettes. Namely
$U_{\rm hop}=\underbrace{\prod_{\begin{subarray}{c}{\rm down-spin}\\ {\rm P-plaquette}\end{subarray}}U_{\rm hop}^{\rm plaq}\prod_{\begin{subarray}{c}{\rm up-spin}\\ {\rm Q-plaquette}\end{subarray}}U_{\rm hop}^{\rm plaq}}_{=:\exp(-itB)}\underbrace{\prod_{\begin{subarray}{c}{\rm down-spin}\\ {\rm Q-plaquette}\end{subarray}}U_{\rm hop}^{\rm plaq}\prod_{\begin{subarray}{c}{\rm up-spin}\\ {\rm P-plaquette}\end{subarray}}U_{\rm hop}^{\rm plaq}}_{=:\exp(-itA)}\,$, (S26)
@@
where $i^{\prime}$ and $j^{\prime}$ denote the qubits on the same physical sites as $i$ and $j$, but with opposite spins. For example, if $i=36$, then $i^{\prime}=0$; if $i=7$ then $i^{\prime}=43$. We note that in this expression, the hopping term in the middle is applied on the qubits $i,j$ (that are not linked by an edge in the fermionic encoding), not on the qubits $i^{\prime},j^{\prime}$ that are linked by an edge in the fermionic encoding. Gathering everything, we obtain the following expression for the circuit that implements one Trotter step under the hopping Hamiltonian:
@@
The main physical effect of this Trotterisation scheme is thus to break the full translation invariance to a symmetry under translation by two sites. Note that, for a periodic system with side length $L=4$, even the next-nearest-neighbour terms in Fig. S9 cancel and thus, for a $4\times 4$ system, our scheme implements the simulation of the hopping term exactly, without Trotter error. Furthermore, for arbitrary system sizes, the second-order Floquet Hamiltonian of the hopping Trotterisation (as well as the term that is dominant for strong interactions $U\gg t$) is shown in Fig. S10. It is noteworthy that part of the double-commutator recovers the original hopping Hamiltonian $H_{\mathrm{hop}}=A+B$. That commutator thus does not lead to an undesired term in the Floquet Hamiltonian, but merely renormalises the hopping $\sim(1+\tau^{2})$ which could in principle be accounted for when targeting a specific evolution time (although we have not made use of this fact for the circuits in this work).
@@
Note that this is a geometrically non-local circuit which allows us to prepare the toric code on the $3\times 3$ periodic ancilla lattice using a two-qubit depth of 4. This is the reason that we are not using a scheme based on measurement and feed-forward: While such schemes offer circuit depths that are constant in system size, for this particular system size, the minimal depth of 4 can be achieved unitarily with a smaller number of overall gates  [iqbal_topological_2024, fossfeig2023experimentaldemonstrationadvantageadaptive].
@@
this time with $\delta=-2/3$ exactly. The discussion of the bilayer perturbation theory is similar to the one in [schlomer_local_2024], although in that reference an exchange coupling between two layers of tJt-J-models was considered, whereas here we allow for the perturbation $V$ to create virtual states with doubly-occupied sites and only turn on the on-site interaction during the adiabatic evolution later on.
@@
The objective function to maximize is the overlap between the target state and the state generated by the circuit. The optimization over the circuit parameters is carried out through gradient descent over the infidelity. Specifically, we utilize the quasi-second order gradient descent method, L-BFGS algorithm, from SciPy [2020SciPy-NMeth]. With the 10-layer circuit, we obtain the fidelity $\mathcal{F}=0.99478$ with the best circuit realization from 30 random initializations. Notice each general $U(1)$ symmetry-preserving gate requires 3 elementary entangling 2-qubit gates, e.g., ZZPhase on our hardware. Due to a cancellation at the first layer, the total depth in terms of elemantary entangling 2-qubit gate of the XXZ state preparation circuit is 29. The parameters for all $10\times 4=40$ general $U(1)$-preserving gates used in the quantum experiments are given in Listing 7.
@@
To carry out the optimization of finding parameters of the fermionic circuits, here we utilize the Jordan-Wigner transformation for simplicity and efficiency. The Jordan-Wigner ordering we choose follows the qubit indices ordering. We first obtained both the $|s\rangle$ and $|d\rangle$ states with the Jordan-Wigner transformation. Then, we transform the variational fermionic circuit correspondingly according to the Jordan-Wigner ordering into a variational circuit over qubits by inserting fSWAPs. That is for any fermionic gate acting over not adjacent fermionic modes in the Jordan-Wigner ordering, we insert the fSWAPs. For example, the $U^{f}_{(1,3)}=\textrm{fSWAP}_{(1,2)}U_{(2,3)}\textrm{fSWAP}_{(1,2)}$. For a fermionic gate acting over adjacent fermionic modes, we apply no additional transformation. Now the problem reduces back to the standard variational quantum circuit optimization. We optimize the circuit with the quasi-second order gradient descent method, L-BFGS algorithm, from SciPy [2020SciPy-NMeth] over 30 different random initialization and select the best result. The best average fidelity is (2C)/2=0.99446(2-C)/2=0.99446.
@@
We use exact state vector simulation to obtain the exact energy gradient of all parameters. The optimization is carried out with the quasi-second order gradient descent method, L-BFGS algorithm, from SciPy [2020SciPy-NMeth] over 30 random initializations.